# Img-Diff: Contrastive Data Synthesis for Multimodal Large Language Models

### v2

@@ 4.4. Results on Image-Editing-Request
As shown in Table 2, both LLaVA-1.5-7B and MGM-7B originally show SOTA performance on the Image-Edit-Request benchmark. Upon incorporating our “object replacement” data for further finetuning, the performance of both LLaVA-1.5-7B and MGM-7B improves even more, achieving new SOTA scores. This suggests that our dataset enhances MLLMs’ abilities to recognize similarities and dissimilarities in image pairs and enables them to describe differences more accurately.
4.5. Results on MLLM Benchmarks
Aside from evaluations related to image difference discrimination, we also assess the performance of our “object replacement” data in enhancing the comprehensive abilities of MLLMs. We test our dataset using commonly used MLLM benchmarks, including VQAv2(Goyal et al., 2017) and GQA(Hudson and Manning, 2019) for assessing the comprehensive VQA capabilities; MMBench(Liu et al., 2023a), MM-Vet(Yu et al., 2023), ScienceQA(Lu et al., 2022), and SEED-Bench(Li et al., 2023c) for testing perceptual and reasoning abilities; and POPE(Li et al., 2023a) for evaluating fine-grained object localization abilities. Table 3 presents the results on these MLLM benchmarks, with the $\triangle$ metric indicating the percentage improvement averaged across them.
Based on Table 3, it can be observed that after finetuning with our dataset, the performance of LLaVA-1.5-7B shows a comprehensive improvement, with an average increase of 3.06% across all benchmarks. For MGM-7B, the improvements brought by our dataset are not as pronounced as those observed with LLaVA-1.5-7B, but it still achieves an average increase of 1.28%. These improvements indicate that the finetuned MLLMs not only enhance the ability to discern differences but also improve overall visual capabilities, thereby better addressing VQA tasks.
4.6. Ablation Studies
@@
Model (3) increases both the CS threshold and the CITM threshold compared to Model (2). Raising the CS threshold implies a greater filtering strength for similar captions, which means that if the two objects corresponding to the same bounding box coordinate in an image pair are similar, the bounding box will be filtered out. Additionally, increasing the CITM threshold aims to enhance the alignment between the captions and the objects being described. After raising both the CS and CITM thresholds, the model’s performance shows a significant improvement.
@@
Based on Table 4, it can be concluded that the stronger the filtering intensity, the better our dataset’s effectiveness. However, due to the increased filtering intensity resulting in a reduced number of final instances, we choose the settings of Model (3) as our optimal threshold to ensure a sufficient number of generated instances. In our future work, we plan to expand the data sources to generate more pairs of similar images and then evaluate the effects of data obtained with higher filtering intensity.
5. Evaluation of Data Quality and Diversity
5.1. Data Quality
To assess the quality of our Img-Diff dataset, we randomly select 1,000 instances of “object replacement” data and employ multiple professional dataset annotators to evaluate the samples based on three metrics. The final scores are determined through a voting process. Specifically, the first metric is “Bounding Box Difference”, which evaluates whether there are differences between the two sub-images of the same bounding box coordinate. If the objects are different, we score it as “high”; if the objects are the same but their features (such as color, shape, etc.) are different, we score it as “medium”; if the objects are the same and their features are similar, we score it as “low”. The second metric is “Content Caption Accuracy”, which evaluates whether the captions generated by Stage 1 of the Difference Captions Generator accurately describe the sub-images. If both captions of an image pair are correct, we score it as “high”; if the captions identify the objects but incorrectly describe their features, we score it as “medium”; if the captions incorrectly identify the objects, we score it as “low”. The third metric is “Difference Caption Accuracy”, which evaluates whether the final difference captions accurately describe the object differences between the image pairs. If the description is accurate, we score it as “high”; if the object recognition is correct but the feature description is incorrect, we score it as “medium”; if the object recognition is incorrect, we score it as “low”. The results are shown in Figure 6.
Figure 6. Quality of the “object replacement” data.
[…]
Quality of the “object replacement” data with manual evaluation.
Based on Figure 6, our dataset demonstrates a high level of quality. For the “Bounding Box Difference” metric, only 4.5% of the instances are classified as “low”, and nearly 80% of instances exhibit significant object differences between their two sub-images. In terms of “Content Caption Accuracy”, 80.1% of sub-image pairs are described accurately, indicating that using an MLLM for labeling is effective and that our filtering strategy is also functioning well. For the “Difference Caption Accuracy” metric, over 70% of the difference descriptions are completely accurate, with 21.8% of the samples having errors solely in feature labeling while still maintaining correct descriptions of object differences, which underscores the effectiveness of our difference caption generation strategy.
5.2. Data Diversity
The effectiveness of a dataset is closely related to the diversity of its data. By analyzing the valid object nouns included in the captions of our image pairs, we assess the diversity of our “object replacement” data. Specifically, we count the total number of object categories covered, the total number of unique “object replacement pairs”, and the frequency of occurrences for each object name. The statistical data is presented in Table 5.
Table 5. Diversity of the “object replacement” data.
Categories Pairs Total Avg Avgobj365
1,203 3,680 25,288 21.02 36.07
As indicated in Table 5, the number of object categories covered by our dataset is 1,203, encompassing most objects in real life. The term “object replacement pair” refers to the combination of the replaced and replacing object names. The total number of unique “object replacement pairs” covered by our dataset is 3,680, with an average of 3.4 occurrences per pair. The total number of occurrences for object names in our dataset is 25,288, with an average of 21.02 occurrences per name.
To evaluate the coverage of common object categories in our dataset, we analyze the occurrences of the object names from the Object365(Shao et al., 2019) dataset in our own dataset. The results show that each object name from the Object365 dataset appears an average of 36.07 times in our dataset, totaling 13,164 occurrences, which accounts for approximately 52.06% of the total occurrences of object names. This indicates that our data instances include a substantial number of common object names, ensuring a high frequency of common objects. Additionally, less common object names make up nearly half of our dataset, demonstrating the broad coverage and high diversity of our dataset.
6. The “Object Removal” Exploration
In addition to object replacement, determining the presence or absence of objects is also crucial. We reform the “object replacement” workflow by integrating an object erasing process, which generates an extended dataset that focuses on object removal. This new data prompts MLLMs to analyze which image in an image pair contains a specific object.
By incorporating this “object removal” data into the “object replacement” dataset, we observed an improvement in performance on eight MLLM benchmarks for LLaVA-1.5-7B, increasing from 3.06% to 3.91%. In contrast, the improvement for MGM-7B slightly decreased from 1.28% to 1.1%. One possible reason for this is that the original training data for MGM-7B has already included a substantial amount of new object existence data compared to LLaVA-1.5.
Based on this exploration, we suggest including “object removal” data as a valuable option to enhance MLLMs, particularly when the finetuning data for MLLMs is limited and lacks sufficient examples related to object presence. More details regarding the ”object removal” data are presented in Appendix A.
7. Conclusion
In this paper, we draw inspiration from recent advances in contrastive learning and image difference captioning to propose a novel method of contrastive data synthesis, creating a high-quality dataset called “Img-Diff” that focuses on describing object differences. Specifically, we generate many pairs of similar images where the main focus is on object replacement. Then, we use the proposed Difference Area Generator and Difference Captions Generator to generate difference captions for specific regions and form question-answer pairs. In contrast to previous image difference datasets, our dataset focuses exclusively on specific regions inside the images, circumventing the issue where the descriptive information of differences does not fully capture the differences between image pairs, thereby enhancing the accuracy of the data. Afterwards, we finetune LLaVA-1.5-7B and MGM-7B using our relatively small-scale dataset, yielding high-performance scores on par with SOTA models trained with much larger-scale datasets in image difference tasks, and comprehensive performance improvements in numerous widely recognized MLLM benchmarks. These results confirm the effectiveness of our dataset to improve the ability of MLLMs in recognizing detailed differences between images.
In a nutshell, we provide a series of insights about the construction of high-quality image difference datasets, showing great potential to effectively and efficiently enhance MLLMs via contrastive data-centric approaches. We meticulously detail the process of creating our Img-Diff dataset and demonstrate the enhanced performance of MLLMs finetuned with our data on both image difference benchmarks and MLLM benchmarks. With this work, we hope it can catalyze further investigation into the realm of image difference datasets and the fine-grained image recognition capabilities of MLLMs.
Appendix
Figure 7. An overview of the generation steps for “object removal” data.
[…]
An overview of the generation process for “object removal” data.
Table 6. Performance comparison on 8 MLLM benchmarks (including “object removal” data).
Model VQAv2 GQA POPE MMB MMBCN $M$-Vet SQAI SEED $\triangle$
LLaVA-7B 78.5 62.0 85.9 64.3 58.3 30.5 66.8 58.6 -
[…]
MGM-7B 80.4 62.6 86.0 69.3 58.9 40.8 70.6 63.5 -
[…]
Table 7. Results on image difference benchmarks (including “object removal” data).
Model MMVP Spot-the-Diff Image-Edit-Request
BLEU METEOR CIDEr-D ROUGE-L BLEU METEOR CIDEr-D ROUGE-L
LLaVA-1.5-7B 24.0 8.5 12.0 38.3 30.1 15.1 17.8 60.6 45.2
LLaVA-1.5-7B+RP 27.3 9.7 13.0 43.2 30.8 16.2 19.5 60.9 46.7
LLaVA-1.5-7B+RP+RM 28.7 9.8 13.0 46.5 31.5 16.8 18.6 63.9 45.7
MGM-7B 40.0 9.9 12.0 46.3 31.5 16.5 17.7 66.8 44.8
MGM-7B+RP 50.7 10.8 13.1 53.5 33.0 16.6 18.2 68.1 45.7
MGM-7B+RP+RM 37.3 9.1 12.0 43.5 31.1 16.1 17.5 64.1 45.5
Appendix A The “Object Removal” Exploration
A.1. Overview
On the main page, we generate many pairs of similar images focusing on object replacement. Their bounding box regions generally contain objects. However, the ability to determine the object presence is also crucial. Thus, we generate another set of pairs of similar images where the difference lies in the presence or absence of objects, to enhance the model’s ability to determine object presence. We refer to these image pairs as “exist-absent pairs” and the data as “object removal” data.
A.2. Generation Process
A.2.1. Workflow
”Object removal” involves erasing a specific object from an image and then merging the edited image with the original to form an exist-absent pair. The detailed workflow is as follows: first, FastSAM is used to segment the image, which provides a set of bounding boxes and masks. Next, the Image Similarity Filter is applied to filter the bounding boxes and accompanying masks, keeping only those that contain objects. Then, we use the text-to-image generative model SDXL-turbo(Sauer et al., 2023) to inpaint the images with the remaining masks, erasing specific objects from the images and generating exist-absent pairs. Afterward, we use an MLLM to describe the removed object for each exist-absent pair, and a filter is employed to verify the accuracy of the description. Finally, we draw red boxes on images based on the bounding box information and then the object descriptions are converted into multiple-choice questions, such as: “which image has the object related to ‘DESCRIPTION’ within the red bounding box? A. the left image B. the right image.” Here, DESCRIPTION refers to the description of the erased objects. After all processing and filtering, we obtain 5,773 pieces of “object removal” data. The general framework is shown in Figure 7.
A.2.2. Image Similarity Filter.
The structure of the current Image Similarity Filter is identical to the one shown in Figure 3. In the current process, the function of the Image Similarity Filter is to filter out bounding boxes whose corresponding areas do not contain objects. For each image, we need its corresponding image in the image pair generated in Section 3.2 to determine whether its bounding box regions contain objects. Since the image pairs are generated by replacing objects, the difference areas between the two images are highly likely to be the regions containing valid objects. Therefore, for each bounding box, we crop the sub-images from image A (the current image) and image B (the other image in the pair), and then calculate the similarity of these two sub-images. When the similarity is below 0.9, we consider these two sub-images to be different, indicating that the bounding box region contains an object.
A.2.3. Erase objects.
We use the model SDXL-turbo(Sauer et al., 2023) to erase objects based on the masks obtained during segmentation. The prompt is “background, nothing, 8k.” After inpainting, the object in the masked regions is erased, while the rest of the image remains unchanged. Hence, we obtain exist-absent pairs.
A.2.4. MLLM Captioning.
We use the MLLM LLaVA-NEXT to generate descriptions for the erased objects. Specifically, we provide the MLLM with the bounding box coordinate and ask it to describe the corresponding area in the original image. Subsequently, we crop the exist-absent pairs based on the bounding box information and then use the Image-Text Matching Filter to assess the matching degree between the sub-images and descriptions. If the score between the sub-image containing objects and its description is greater than 0.35, and the score between the sub-image not containing objects and its description is less than 0.2, we consider the description to be accurate and the exist-absent pair to be valid.
A.3. Evaluation
We merge the “object removal” data with the “object replacement” data, making our dataset focus on object changes as well as object presence. To test the performance changes of LLaVA-1.5-7B and MGM-7B after adding “object removal” data, we incorporate the combined data into the original training datasets of these two MLLMs for visual instruction tuning and evaluate the finetuned MLLMs on image difference benchmarks and MLLM benchmarks, just like the main page.
In the tables, “RM” represents “object removal” data.
A.3.1. Results on MLLM benchmarks
Table 6 shows the performance of LLaVA-1.5-7B and MGM-7B finetuned with additional “object removal” data on commonly used MLLM benchmarks. With the assistance of “object removal” data, LLaVA-1.5-7B achieves further improvements across various benchmarks compared to the model that only includes “object replacement” data, with an average increase of 3.91%. However, after adding “object removal” data, MGM-7B experiences a slight decline in performance, with an average increase of only 1.10% across all the benchmarks. Nevertheless, it still shows score improvements on the POPE benchmark and the MMBench benchmarks.
A.3.2. Results on image difference benchmarks
Table 7 shows the performance of LLaVA-1.5-7B and MGM-7B finetuned with our “object removal” data on image difference benchmarks. It can be observed that with “object removal” data, LLaVA-1.5-7B’s scores improve on the MMVP benchmark and the Spot-the-Diff benchmark, while its scores fluctuate on the Image-Edit-Request benchmark. However, MGM-7B experiences varying degrees of score decreases across all three image difference benchmarks.
A.3.3. Analysis
The benchmark results indicate that the “object removal” data has a comprehensive positive effect on LLaVA-1.5-7B. However, upon finetuning with the “object removal” data, the performance of MGM-7B shows fluctuations. We speculate that the limited improvement for MGM might be due to the broad scope of MGM’s original data, which may already encompass the new information introduced by our “object removal” dataset. In conclusion, adding “object removal” data does not necessarily lead to a significant performance improvement for MLLMs. However, when finetuning data is limited or does not include data related to the presence or absence of objects, “object removal” data can serve as an option for data augmentation.
Appendix B Additional Details of Experiments
B.1. Preprocessing of image pairs before inputting into MLLMs
The MLLMs selected in our paper (LLaVA-1.5, LLaVA-NEXT and MGM) only support single-image input. Therefore, our image pairs need to be horizontally concatenated before being fed into MLLMs’ image encoder. Specifically, we horizontally concatenate the image pair and add a vertical black dividing line, 20 pixels wide, between the images.
B.2. Training Process for MLLMs
The training process for advanced MLLMs, such as LLaVA-1.5 and MGM, typically involves two stages: pretraining and finetuning. During the pretraining phase, the MLLMs keep the backbone LLM and vision encoder frozen and zero-initialize the learnable projector which is usually an MLP and used for semantic mapping and cross-modality alignment. Only the projector is trained using the pretraining dataset. In the finetuning phase, the MLLMs unfreeze the backbone LLM and finetune both the backbone LLM and the learnable projector using the visual instruction tuning dataset. Specifically, the pretraining dataset is usually an image captioning dataset, while the visual instruction tuning dataset typically consists of VQA datasets for various tasks. Thus, our Img-Diff dataset is integrated into the visual instruction tuning dataset during the finetuning phase and used together with the original dataset to finetune the MLLMs.
B.3. Model Selection
B.3.1. Overview
@@
B.3.2. Selection of the Semantic Segmentation Model
In our project, we need to use a semantic segmentation model to identify regions containing objects in images. To ensure a diverse range of object categories is covered, we opt for models like SAM(Kirillov et al., 2023) instead of traditional semantic segmentation models. Furthermore, to reduce time consumption, we select FastSAM, one of the most efficient and effective models within SAM-like category, as our segmentation model.
B.3.3. Model Size
Considering the device limitation and time consumption, our paper utilizes the LLM Vicuna-1.5-13B for object name replacement in the image pairs generation process. For semantic segmentation in the Difference Area Generator, the FastSAM-x model is employed. For the CLIP model, we choose “clip-vit-base-patch32”, and for the BLIP model, we select “blip-itm-large-coco”. In the Difference Captions Generator, we use the MLLM LLaVA-NEXT-13B to generate content captions and difference captions. These models are interchangeable. When resources allow, researchers can substitute them with higher-performance models to achieve datasets with improved performance.
B.4. Filtering Thresholds
During the generation process of “object replacement” data, we employ multiple filtering operations. In this subsection, we outline the filtering thresholds we use.
@@ B.4. Filtering Thresholds
Setting the filtering intensity too high may lead to a reduced number of remaining samples. To ensure that the dataset still has enough samples after filtering, we outline adjustable thresholds as described above. As mentioned in Section 4.6, higher filtering intensity typically results in better model performance. Therefore, researchers may consider expanding the data sources and increasing the filtering intensity to improve dataset performance.
Appendix C Examples
Figure 8. Examples of “object replacement” data, including the image pair and the text content in JSON format.
@@ Appendix C Examples
Examples of “object replacement” data, including the image pair and the text content in JSON format.
Figure 9. An example of “object removal” data, including the image pair and the text content in JSON format.
@@
An example of “object removal” data, including the image pair and the text content in JSON format.

### v3

@@ 14 Ablation Studies
Model (3) increases both the CS threshold and the CITM threshold compared to Model (2). Raising the CS threshold implies a greater filtering strength for similar captions, which means that if the two objects corresponding to the same bounding box coordinate in an image pair are similar, the bounding box will be filtered out. As for the CITM threshold, increasing the CITM threshold aims to enhance the alignment between the captions and the objects being described. After raising both the CS and CITM thresholds, the model’s performance shows a significant improvement.
@@
Based on Table 14, it can be concluded that the stronger the filtering intensity, the better our dataset’s effectiveness. However, due to the increased filtering intensity resulting in a reduced number of final instances, we choose the settings of Model (3) as our optimal threshold to ensure a sufficient number of generated instances. In our future work, we will expand the data sources to generate more pairs of similar images and then evaluate the effects of data obtained with higher filtering intensity.
15 Additional Details of Experiments
15.1 Preprocessing of image pairs before inputting into MLLMs during training and inference
The MLLMs selected in our paper (LLaVA-1.5, MGM, InternVL2) only support single-image input. Therefore, our image pairs need to be horizontally concatenated before being fed into MLLMs’ image encoder. Specifically, we horizontally concatenate the images in pairs and add a vertical black dividing line, 20 pixels wide, between the images.
15.2 Training Process for MLLMs
The training process for advanced MLLMs, including LLaVA-1.5, MGM and InternVL2, typically involves two stages: the pre-training stage and fine-tuning stage. During the pre-training stage, the MLLMs keep the backbone LLM and the vision encoder frozen and zero-initialize the learnable projector which is used for semantic mapping and cross-modality alignment. Only the projector is trained using the pre-training dataset. In the fine-tuning stage, we unfreeze the backbone LLM and fine-tune both the backbone LLM and the learnable projector using the visual instruction tuning dataset. Specifically, the pre-training dataset is usually an image captioning dataset, while the visual instruction tuning dataset typically consists of VQA datasets for various tasks. Thus, the Img-Diff dataset is integrated into the visual instruction tuning dataset during the fine-tuning stage and used together with the original dataset to fine-tune the MLLMs.
15.3 Model Selection
15.3.1 Overview
@@
15.3.2 Selection of the Semantic Segmentation Model
In our project, we need to use a semantic segmentation model to identify regions containing objects in images. To ensure a diverse range of object categories is covered, we opt for models like SAM [31] instead of traditional semantic segmentation models. Furthermore, to reduce time consumption, we select FastSAM, one of the most efficient and effective models within the SAM-like category, as our segmentation model.
15.3.3 Model Size
Considering the device limitation and time consumption, our paper utilizes the LLM Vicuna-1.5-13B [13] for object name replacement in the image pairs generation process. For semantic segmentation in the Difference Area Generator, the FastSAM-x model is employed. For the CLIP model, we choose “clip-vit-base-patch32”, and for the BLIP model, we select “blip-itm-large-coco”. In the Difference Captions Generator, we use the MLLM LLaVA-NEXT-13B to generate content captions and difference captions. These models are interchangeable. When resources allow, researchers can substitute them with higher-performance models to achieve datasets with improved performance.
15.4 Filtering Thresholds
During the generation process of “object replacement” data, we employ multiple filtering operations. In this subsection, we will outline the filtering thresholds we use.
@@ 15.4 Filtering Thresholds
Setting the filtering intensity too high may lead to a reduced number of remaining samples. To ensure that the dataset still has enough samples after filtering, we outline adjustable thresholds as described above. As mentioned in Section 14, higher filtering intensity typically results in better model performance. Therefore, researchers may consider expanding the data sources and increasing the filtering intensity to improve dataset performance.
15.5 Resource and Time Consumption
With four NVIDIA A100 GPUs, it took 4.5 days to synthesize 118K high-quality image pairs. The subsequent filtering and description-generating processes took approximately two days in total.
Figure 7: Three “object removal” examples.
Figure 8: An overview of the generation steps for “object removal” data.
16 The “Object Removal” Exploration
16.1 Overview
On the main page, we generate pairs of similar images focusing on object replacement. Their bounding box regions generally contain objects. However, the ability to determine the object’s presence is also crucial. Thus, we generate another set of image pairs where the difference lies in the presence or absence of objects, to enhance the model’s ability to determine object presence. We refer to these image pairs as “exist-absent pairs” and the data as “object removal” data.
16.2 Generation Process
16.2.1 Workflow
“Object removal” involves erasing a specific object from an image and then merging the edited image with the original to form an exist-absent pair. The detailed workflow is as follows: first, FastSAM is used to segment the image, which provides a set of bounding boxes and masks. Next, an Image Similarity Filter is applied to filter the bounding boxes and accompanying masks, keeping only those that contain objects. Then, we use the text-to-image generative model SDXL-turbo[57] to inpaint the images with the remaining masks, erasing specific objects from the images and generating exist-absent pairs. Next, we use an MLLM to describe the removed object for each exist-absent pair, and a filter is employed to verify the accuracy of the description. Finally, we draw red boxes on images based on the bounding box information, and then the object descriptions are converted into multiple-choice questions, such as: “which image has the object related to ‘DESCRIPTION’ within the red bounding box? A. the left image B. the right image.” Here, DESCRIPTION refers to the description of the erased objects. After all processing and filtering, we obtain 5,773 pieces of “object removal” data. The general framework is shown in Figure 8.
16.2.2 Image Similarity Filter
In the current process, the function of the Image Similarity Filter is to filter out the bounding box regions that do not contain objects. For each image, we need its corresponding image in the image pair generated in Section 3.2 to determine whether its bounding box regions contain objects. Since the image pairs are generated by replacing objects, the difference areas between the two images are highly likely to be the regions containing valid objects. Therefore, for each bounding box, we crop the sub-images from image A (the current image) and image B (the other image in the pair), and then calculate the similarity of these two sub-images. When the similarity is below 0.9, we consider these two sub-images to be different, indicating that the bounding box region contains an object.
16.2.3 Erase Objects
We use the generative model SDXL-turbo[57] to erase objects based on the masks obtained during segmentation. The prompt is “background, nothing, 8k.” After inpainting, the object in the masked regions is erased, while the rest of the image remains unchanged. Hence, we obtain exist-absent pairs.
16.2.4 MLLM Captioning
We use the MLLM LLaVA-NEXT to generate descriptions for the erased objects. Specifically, we provide the MLLM with the bounding box coordinates and ask it to describe the corresponding area in the original image. Subsequently, we crop the exist-absent pairs based on the bounding box information and then use an Image-Text Matching Filter to assess the matching degree between the sub-images and the descriptions. If the matching score between the sub-image containing objects and its description is greater than 0.35, and the matching score between the sub-image not containing objects and its description is less than 0.2, we consider the description to be accurate and the exist-absent pair to be valid.
Table 15: Performance comparison on 8 MLLM benchmarks (including “object removal” data).
Model VQAv2 GQA POPE MMB MMBCN
LLaVA-7B 78.5 62.0 85.9 64.3 58.3
LLaVA-7B + RP 79.3 62.8 86.4 66.1 59.8
LLaVA-7B + RP + RM 79.2 62.9 86.8 67.9 61.3
Model $M$-Vet SQAI SEED $\triangle$ MMVP
LLaVA-7B 30.5 66.8 58.6 - 24.0
@@ 16.2.4 MLLM Captioning
Table 16: Results on image difference benchmarks (including “object removal” data).
Model Spot-the-Diff
BLEU METEOR CIDEr-D ROUGE-L
LLaVA-1.5-7B 8.5 12.0 38.3 30.1
LLaVA-1.5-7B +RP 9.7 13.0 43.2 30.8
LLaVA-1.5-7B +RP +RM 9.8 13.0 46.5 31.5
Model Image-Edit-Request
BLEU METEOR CIDEr-D ROUGE-L
LLaVA-1.5-7B 15.1 17.8 60.6 45.2
LLaVA-1.5-7B +RP 16.2 19.5 60.9 46.7
LLaVA-1.5-7B +RP +RM 16.8 18.6 63.9 45.7
16.3 Evaluation
We merge the “object removal” data with the “object replacement” data, making our dataset focus on both object changes and object presence. To test the performance changes of LLaVA-1.5-7B after adding “object removal” data, we incorporate this combined data into the original visual instruction tuning dataset of the MLLM and conduct fine-tuning. Then, we evaluate the fine-tuned model on image difference benchmarks and eight MLLM benchmarks, similar to what is presented on the main page.
In the tables, “RM” represents “object removal” data.
16.3.1 Results on MLLM Benchmarks
Table 15 shows the performance of LLaVA-1.5-7B finetuned with additional “object removal” data on commonly used MLLM benchmarks. With the assistance of “object removal” data, LLaVA-1.5-7B achieves further improvements across various benchmarks compared to the model that only uses “object replacement” data, with an average increase of 3.91%.
16.3.2 Results on Image Difference Benchmarks
Table 16 shows the performance of LLaVA-1.5-7B fine-tuned with our “object removal” data on image difference benchmarks. With “object removal” data, LLaVA-1.5-7B shows further improvements in its performance on both the MMVP benchmark and the Spot-the-Diff benchmark, surpassing the results achieved with “object replacement” data alone. Besides, its scores fluctuate on the Image-Edit-Request benchmark.
16.3.3 Further Analysis
The results indicate that the “object removal” data has a comprehensive positive impact on LLaVA-1.5-7B, leading to performance improvements in both MLLM benchmarks and image difference benchmarks. However, during our analysis of sample quality, we notice that some of the generated “object removal” samples exhibit subpar image quality, with certain image pairs showing inadequate object removal effects. In light of this, employing a more robust inpainting model or applying additional filters to enhance the quality of these image pairs could further optimize the performance of this dataset.
17 Examples
Figure 9: Examples of “object replacement” data, including the image pair and the text content in JSON format.
Figure 10: An example of “object removal” data, including the image pair and the text content in JSON format.